OpenStudy (anonymous):

A quadratic equation is shown below: 3x2 - 16x + 2 = 0 Part A: Describe the solution(s) to the equation by just determining the discriminant. Show your work. (3 points) Part B: Solve 9x2 + 3x - 2 = 0 by using an appropriate method. Show the steps of your work, and explain why you chose the method used. (4 points) Part C: Solve 3x2 - 10x + 2 = 0 by using a method different from the one you used in Part B. Show the steps of your work. (3 points)

OpenStudy (anonymous):

@YanaSidlinskiy

OpenStudy (yanasidlinskiy):

Formula for quadratics would be: \[ax^2+bx+c\]

OpenStudy (anonymous):

oops sorry the equation is 3x^2 - 16x + 2 = 0

OpenStudy (yanasidlinskiy):

Can you post it up again? Because you have different equations..In the parts^

OpenStudy (anonymous):

In the equation: \(ax^2+bx+c=0\)' \(discriminant\ (D)=b^2-4ac\), to solve for the zeroes/solutions, you can use the quadratic formula \(\Large x = \frac {-b \pm \sqrt {b^2 - 4ac}}{2a}\)